The average weight of 80 randomly


selected sacks of rice is 45.54 kilos with a


standard deviation of 17 kilos. Test the


hypothesis at a 0.01 level of significance that


the true mean weight is less than 49 kilos. 2. The average length of time for students to


have their subjects controlled is 40 minutes. A


new controlling procedure using modern


computing machines is being tried. If a random


sample of 15 students has an average


controlling time of 25 minutes with a standard


deviation of 12.9 minutes under the new


system, test the hypothesis that the average


length of time to control student’s subjects is


less than 40 minutes. Use a level of


significance of 0.10 and assume the


population of controlling times to be normally


distributed.

1
Expert's answer
2022-05-18T07:00:40-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge49"

"H_a:\mu<49"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=79" degrees of freedom, and the critical value for a left-tailed test is "t_c =-2.374482."The rejection region for this left-tailed test is "R = \{t:t< -2.374482\}."

The t-statistic is computed as follows:


"t=\dfrac{45.54-49}{17/\sqrt{80}}=-1.8204"

Since it is observed that "t = -1.8204> -2.374482=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed "df=79" degrees of freedom, "t=-1.8204" is "p=0.036243," and since "p=0.036243>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 80, at the "\alpha = 0.01" significance level.


2. The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge40"

"H_a:\mu<40"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=14" degrees of freedom, and the critical value for a left-tailed test is "t_c =-1.34503."The rejection region for this left-tailed test is "R = \{t:t< -1.34503\}."

The t-statistic is computed as follows:


"t=\dfrac{25-40}{12.9/\sqrt{15}}=-4.503469"

Since it is observed that "t = -4.503469<-1.34503=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed "df=14" degrees of freedom, "t=-4.503469" is "p= 0.000248," and since "p= 0.000248<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 40, at the "\alpha = 0.10" significance level.